OpenStudy (anonymous):

I need help it says solve for w when the formula is p=4+2w I dont know what to do with the p

OpenStudy (anonymous):

first, get the w term by itself. So move 4 over to the left hand side. Do you know how to do this?

OpenStudy (anonymous):



OpenStudy (anonymous):

that does not explain the steps of the problem....

OpenStudy (anonymous):

what I did I divide 2 on both sides and my answer I got is w=4/2

OpenStudy (anonymous):

the first step: when you have p=4+2w, you first move the 4 over to the left hand side. you do this by subtracting. So your equation become p-4=2w. The next step to getting w by itself is to divide both sides by 2. So we end up with p/2 - 2=w.

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

yes know it does thank you

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

could you help me in 2 more problems

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

ok a=s squared solve for s and the other ones is 3.14r squared solve for r

OpenStudy (anonymous):

s = \[\sqrt{a}\] this is because \[\sqrt{s^{2}}\]=s, and what you do to one side you must do to the other.

OpenStudy (anonymous):

and the other question is just \[3.14r^{2}\]? is it set equal to anything?

OpenStudy (anonymous):

no it just says a=3.14r squared

OpenStudy (anonymous):

ok, so first divide both sides by 3.14

OpenStudy (anonymous):

then you're left with \[\frac{ a }{ 3.14 } = ^{2}\]

OpenStudy (anonymous):

\[r^{2} *\]

OpenStudy (anonymous):

so then we do the same thing as we did in the last problem...what you do to one side you must do to the other

OpenStudy (anonymous):

so \[r = \sqrt{\frac{ a }{ 3.14 }}\]

OpenStudy (anonymous):

ok thank you so munch you made it easier to me to understand I wish you were my teacher instead thank you so much

OpenStudy (anonymous):

Aw well I'm glad I helped!! let me know if you have any other questions

OpenStudy (anonymous):

oK THANK YOU :)